OpenStudy (anonymous):

if one person can mow the lawn in 4 hours and another can mow the lawn in 5 hours how long would it take them if the did it together? Please give me an equation so I can figure this out

OpenStudy (anonymous):

9/20

OpenStudy (anonymous):

one rate is \[\frac{1}{4}\] other rate is \[\frac{1}{5}\] combined rate is \[\frac{4+5}{4\times 5}\] and you wan to solve \[\frac{4+5}{4\times 5}T=1\] for T, get \[\frac{4\times 5}{4+5}=\frac{20}{9}\]

OpenStudy (anonymous):

sorry i had it upside down. it is \[\frac{4\times 5}{4+5}\]

OpenStudy (anonymous):

nevermind I got it